OpenStudy (anonymous):

Find the perimeter of a rectangle with l = 0.85 m and w = 0.4 m. A. 1.25 m B. 2.5 m C. 0.34 m D. 0.68 m

OpenStudy (anonymous):

@FluffyYumYum

OpenStudy (anonymous):

Any pics?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Well just do 0.85+0.4+0.4=?

OpenStudy (anonymous):

xD It's easier with pics

OpenStudy (anonymous):

0.93

OpenStudy (anonymous):

wait woops

OpenStudy (anonymous):

I was like... Huh?

OpenStudy (anonymous):

3.3

OpenStudy (anonymous):

Nevermind just do one 0.4 because then it adds up to 1.25

OpenStudy (anonymous):

0.85+0.4=1.25

OpenStudy (anonymous):

And therefor it's A

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

oh well it is right xD

OpenStudy (anonymous):

need any more help?

OpenStudy (anonymous):

just did all the math lol with a calculator

OpenStudy (anonymous):

yea if I can find some I need help with

OpenStudy (anonymous):

Okay just tag me with it and i'll help you

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Medal? :P

OpenStudy (anonymous):

Yayyyyy

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24 @FluffyYumYum

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I'd first add 18+24=?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

42

OpenStudy (anonymous):

stupid computer

OpenStudy (anonymous):

If we did A which is 18*24 that's 432 and thats to big so cross that off

OpenStudy (anonymous):

already done that like 5 billion times to it lol

OpenStudy (anonymous):

If we do B 2* 18+24= 60. and we already figured out thats its 42 so its to high by just a little bit

OpenStudy (anonymous):

If we do C which is 2* 18+2* 24 we get 84 and thats to high xD

OpenStudy (anonymous):

so its D

OpenStudy (anonymous):

D which is 16+24=40 and thats the closest

OpenStudy (anonymous):

So i'd go with B

OpenStudy (anonymous):

D*

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

xD Stupid fingers

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

2 people want to help me sooo brb I'll tag you there

OpenStudy (anonymous):

:P Stupid fingers but smart brain :P

OpenStudy (anonymous):

I mean want me to help them

OpenStudy (anonymous):

Okay